OpenStudy (anonymous):

How do you go about solving this integration where you have (x^2)dy/dx+xy=1 it is called first order linear equation

OpenStudy (anonymous):

x(x y'+y)=1 x y'+y=1/x d/dx(x*y(x))=1/x x*y(x)=log(x)+C y(x)=log(x)/x+C/x

OpenStudy (anonymous):

Thanks that is great except it would be lnx. I saw a different method in the book and was looking for something simpler and yours is simpler. how do I give a medal I am new on this site.

OpenStudy (anonymous):

asked someone think I have given you a medal